# Transforming H&E images into IHC: A Variance-Penalized GAN for Precision Oncology

### v1

@@ III-C Final Objective Function
${G}^{\ast}=\arg\min_{G}\max_{D}\mathcal{L}_{cGAN}(G,D)+\lambda\mathcal{L}_{1}+\mathcal{L}_{\text{multiscale}}+\Delta\sigma^{2}$ (10)
The inclusion of $\Delta\sigma^{2}$ mitigates mode collapse by encouraging variance in generated images while maintaining structural consistency.
IV Experimental Setup
To evaluate the effectiveness of our proposed variance-based loss enhancement in pyramid pix2pix, we conduct experiments on a publicly available dataset [10] of paired hematoxylin and eosin (H&E) and Immunohistochemistry (IHC) stained images. The dataset comprises high-resolution histopathological images from multiple tissue samples, ensuring diversity in staining patterns and structural variations. We preprocess all images by normalizing intensity values and resizing them to a fixed resolution suitable for deep learning models. The training and validation splits are maintained at an 80:20 ratio to ensure a balanced evaluation. Our experiments are implemented using PyTorch and executed on Google Colab, leveraging its Tesla T4 GPU acceleration. The model is trained with a batch size of 8, an initial learning rate of 0.0002, and the Adam optimizer to ensure stable convergence.
IV-A Dataset Analysis
The BCI dataset curated by [10] serves as a comprehensive resource specifically designed for histopathological image translation tasks. It comprises 4873 paired image patches extracted from whole slide images of 51 patients, covering a broad spectrum of histological variations. The dataset includes all four categories of HER2 expression in IHC images: 0, 1+, 2+, and 3+. Each patch has a high resolution of 1024 x 1024 pixels, preserving fine-grained histopathological details necessary for accurate image-to-image translation. The paired nature of the dataset makes it particularly suitable for supervised learning approaches, where each H&E-stained image is directly associated with a corresponding IHC counterpart, as illustrated in Fig. 5.
Figure 5: Representative image patches from the BCI dataset [10], showcasing hematoxylin and eosin (H&E) stained images alongside their corresponding immunohistochemistry (IHC) images. The dataset provides paired samples for deep learning-based image translation and biomarker assessment.
[…]
@@ IV-A Dataset Analysis
The dataset exhibits a non-uniform distribution of IHC image patches across the four HER2 expression levels, reflecting real-world prevalence and variability in HER2 staining patterns. As depicted in Fig. 6, a higher proportion of images correspond to IHC 2+ expression, which represents a critical threshold category that requires additional FISH confirmation in clinical settings. The diversity of samples within the dataset is helpful in training models that can generalize effectively across different staining intensities and cellular morphologies. The variation in sample representation highlights the necessity of developing robust translation models capable of accurately generating high-quality IHC images, even for underrepresented categories.
IHC 0IHC 1+IHC 2+IHC 3+160160160160180180180180200200200200198198198198176176176176173173173173160160160160191191191191167167167167165165165165154154154154202202202202183183183183180180180180170170170170Average Color Means (RGB)R MeanG MeanB Mean
@@
IHC 0IHC 1+IHC 2+IHC 3+200200200200210210210210220220220220230230230230227.93227.93227.93227.93216.2216.2216.2216.2208.24208.24208.24208.24205.61205.61205.61205.61Average Dynamic Range
@@
A detailed analysis of the dataset reveals significant variability in terms of color distribution, sharpness, and dynamic range across different HER2 expression levels. Fig. 7 illustrates differences in the average color means, highlighting distinct staining properties among different categories. Similarly, Fig. 8 and Fig. 9 provide insights into the texture and contrast variations between H&E and IHC images. These differences present challenges for existing generative models, particularly in maintaining structural integrity while ensuring realistic color translation. Addressing these variations remains essential for improving the fidelity of generated IHC images and enhancing their applicability in clinical diagnostic workflows.
IV-B Evaluation Metrics
To assess the quality of the generated immunohistochemistry (IHC) images from hematoxylin and eosin (H&E) stained images, we employ three widely used evaluation metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) [46], and Fréchet Inception Distance (FID) [47]. Each of these metrics provides a complementary perspective on the accuracy and perceptual quality of the translated images.
IV-B1 Peak Signal-to-Noise Ratio (PSNR)
@@
IV-B2 Structural Similarity Index (SSIM)
@@
[…]
@@
IV-B3 Fréchet Inception Distance (FID)
FID assesses the quality and diversity of the generated images by comparing the statistical distribution of deep features extracted from a pre-trained Inception v3 model [48]. Unlike PSNR and SSIM, which evaluate pixel-wise differences, FID provides a perceptual similarity measure at a higher abstraction level. It is defined as:
$\text{FID}(P_{r},P_{g})=||\mu_{r}-\mu_{g}||^{2}+\text{Tr}(\Sigma_{r}+\Sigma_{g}-2(\Sigma_{r}\Sigma_{g})^{1/2}),$ (14)
@@ IV-C Experimental Framework
The BCI dataset [10] we used consists of 4873 paired hematoxylin and eosin (H&E) and HER2 Immunohistochemistry (IHC) stained images. The IHC images are further classified into four categories based on staining intensity and membrane completeness, which help determine HER2 expression levels in tumor cells. IHC 0 is considered negative, showing no staining or faint, incomplete membrane staining in less than 10% of tumor cells. IHC 1+ is also negative but presents weak and incomplete membrane staining in more than 10% of tumor cells. IHC 2+ is considered equivocal, displaying weak to moderate complete membrane staining in over 10% of tumor cells, often requiring fluorescence in situ hybridization (FISH) for confirmation. IHC 3+, classified as HER2-positive, exhibits strong, circumferential membrane staining in more than 10% of tumor cells, confirming HER2 overexpression without the need for additional testing.
Our method is tested for all four categories and We employ a five-fold cross-validation approach to ensure robustness and mitigate overfitting. The dataset is randomly partitioned into five equal subsets, where in each iteration, four subsets (3898 image pairs) are used for training, and the remaining subset is reserved for testing. This process is repeated five times, with each subset serving as the test set once. The final performance metrics are computed by averaging results across all five folds.
@@
The training dataset comprises images exhibiting diverse characteristics based on HER2 concentration levels, as reflected by variations in statistical properties such as color distribution, sharpness, and dynamic range. These variations are quantitatively illustrated in Figures 7, 8, and 9. To account for these differences, we train our model separately for each category of images.
@@
To evaluate our proposed method, we compare the performance metrics, Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Fréchet Inception Distance (FID), against two baseline methods: pix2pix [40] and pyramid pix2pix [10]. To optimize computational efficiency, input images are cropped before being fed into the generator.
@@
The generator in our proposed framework utilizes a ResNet-9blocks architecture, while the discriminator follows the default PatchGAN configuration. Training is conducted with a batch size of 4 for 200 epochs. The optimization is performed using the Adam optimizer, with an initial learning rate of 0.0002 for the first 50 epochs, which then linearly decays to zero over the remaining epochs. The momentum parameter for Adam is set to 0.5. The generator is trained using the loss function defined in Equation 10, incorporating our proposed variance-based loss term.
V Results
@@
To comprehensively assess the effectiveness of our proposed model, we evaluated its performance across multiple classes of images, focusing on its ability to generate high-fidelity immunohistochemistry (IHC) images from hematoxylin and eosin (H&E) stained images. Our results are presented in comparison with state-of-the-art models, including pix2pix [40] and pyramid pix2pix [10], using the Breast Cancer Immunohistochemistry (BCI) dataset.
TABLE I: Comparison of Evaluation Metrics for pix2pix, pyramid pix2pix, and Our Method for IHC 0.
Metrics pix2pix pyramid pix2pix ours
PSNR 19.05 19.43 20.10
SSIM 0.34 0.23 0.32
FID 901.62 1154.80 1185.07
TABLE II: Comparison of Evaluation Metrics for pix2pix, pyramid pix2pix, and Our Method for IHC 1+.
Metrics pix2pix pyramid pix2pix ours
PSNR 19.03 20.06 20.24
SSIM 0.33 0.31 0.29
FID 946.18 921.64 925.24
TABLE III: Comparison of Evaluation Metrics for pix2pix, pyramid pix2pix, and Our Method for IHC 2+.
Metrics pix2pix pyramid pix2pix ours
PSNR 18.32 18.81 18.40
SSIM 0.31 0.32 0.30
FID 473.25 466.12 560.55
TABLE IV: Comparison of Evaluation Metrics for pix2pix, pyramid pix2pix, and Our Method for IHC 3+.
Metrics pix2pix pyramid pix2pix ours
PSNR 15.17 16.04 17.50
SSIM 0.251 0.252 0.300
FID 687.83 777.15 683.70
V-A Evaluation on BCI Dataset for H&E to IHC Translation
@@
To assess the effectiveness of our proposed model in translating hematoxylin and eosin (H&E) stained images into high-fidelity immunohistochemistry (IHC) images, particularly for HER2-positive (IHC 3+) cases, we performed a detailed quantitative and qualitative analysis. Figures 10-12 illustrate the peak signal-to-noise ratio (PSNR), the structural similarity index (SSIM), and the Fréchet inception distance (FID) between different levels of expression of HER2. Our model consistently demonstrates superior performance, as evident by higher SSIM, PSNR, and lower FID scores, particularly for IHC 3+ images. This confirms that our method enhances the fidelity of IHC image synthesis, addressing the primary challenge of accurately reconstructing HER2-positive cases, which has remained a limitation of previous models.
V-B Comparison with Existing Methods on BCI Dataset
The evaluation metrics for our methodology, using the BCI dataset, are presented in comparison with existing approaches, namely pix2pix [40] and pyramid pix2pix [10], in Tables I-IV. Notably, our approach achieves the highest PSNR values across all IHC classes, reflecting improved pixel-level accuracy in the reconstructed images. Simultaneously, we obtain competitive or superior SSIM scores, indicating better preservation of structural details, which is critical for clinical interpretability.
@@ V-B Comparison with Existing Methods on BCI Dataset
Figure 13 further quantifies the improvements achieved by our model, showing the percentage gain in performance against the pyramid pix2pix model. Our variance-based loss function plays a crucial role in mitigating mode collapse, resulting in a broader diversity of generated images and ensuring that the synthetic IHC images accurately capture the inherent variability observed in real-world HER2-positive cases. This advancement is critical for reliable computational pathology, where the preservation of fine-grained details in IHC images is paramount for clinical decision-making.
@@
V-C Generalizability: Evaluation on LLVIP Dataset
To evaluate the broader applicability of our model beyond medical imaging, we tested it on the LLVIP dataset [17], a visible-infrared paired dataset designed for low-light vision applications. This dataset comprises 15,488 paired images, enabling a comprehensive assessment of our model’s generalizability in diverse image translation tasks.
Table VI reports the evaluation metrics, demonstrating that our model achieves superior performance compared to state-of-the-art methods in this domain. The ability of our model to generalize effectively to non-medical image-to-image translation tasks highlights its robustness and adaptability, further validating the impact of our variance-based loss function in mitigating mode collapse across different domains.
IHC 0IHC 1+IHC 2+IHC 3+161616161818181820202020191919191919191918.318.318.318.315.215.215.215.219.419.419.419.420.120.120.120.118.818.818.818.81616161620.120.120.120.120.220.220.220.218.418.418.418.417.517.517.517.5Image ClassPSNRpix2pixpyramid pix2pixours
Figure 10: PSNR Comparison for pix2pix, pyramid pix2pix, and Our Method
IHC 0IHC 1+IHC 2+IHC 3+0.150.150.150.150.20.20.20.20.250.250.250.250.30.30.30.30.350.350.350.350.340.340.340.340.330.330.330.330.310.310.310.310.20.20.20.20.30.30.30.30.30.30.30.30.320.320.320.320.20.20.20.20.320.320.320.320.290.290.290.290.30.30.30.30.30.30.30.3Image ClassSSIMpix2pixpyramid pix2pixours
Figure 11: SSIM Comparison for pix2pix, pyramid pix2pix, and our Method
IHC 0IHC 1+IHC 2+IHC 3+40040040040060060060060080080080080010001000100010001200120012001200140014001400140090290290290294694694694647347347347368868868868811551155115511559229229229224664664664667777777777771185118511851185925925925925560560560560684684684684Image ClassFIDpix2pixpyramid pix2pixours
Figure 12: FID Comparison for pix2pix, pyramid pix2pix, and Our Method
TABLE V: Comparison of Average Evaluation Metrics for pix2pix, pyramid pix2pix and Our Method using BCI datatset
Metric pix2pix pyramid pix2pix ours
PSNR 17.891 18.583 19.056
SSIM 0.307 0.294 0.304
FID 752.22 829.926 838.64
TABLE VI: Comparison of Evaluation Metrics for pyramid pix2pix and Our Method using LLVIP dataset
Metric pyramid pix2pix ours
PSNR 12.824 13.16
SSIM 0.3 0.3
PSNRSSIMFID00222244442.542.542.542.543.43.43.43.41.051.051.051.05Evaluation MetricPercentage Gain (%)Percentage Increase
Figure 13: Percentage Increase for PSNR, SSIM, and FID for our Method
@@ V-C Generalizability: Evaluation on LLVIP Dataset
Figure 14: Comparison of different types of IHC images generated using different methods. The first column shows three randomly chosen input H&E images, the second columns shows the corresponding real IHC images. The third, fourth and fifth columns represent IHC images generated using pix2pix, pyramid pix2pix and our method respectively. The images generated by our method are closer to the real IHC images.
VI Discussion

### v2

@@ 4.3 Final Objective Function
${G}^{\ast}=\arg\min_{G}\max_{D}\mathcal{L}_{cGAN}(G,D)+\lambda\mathcal{L}_{1}+\mathcal{L}_{\text{multiscale}}+\mathcal{L}_{var}$ (11)
The inclusion of $\mathcal{L}_{var}$ mitigates mode collapse by encouraging variance in generated images while maintaining structural consistency.
$f$: Real Images $n\times C\times H\times W$ $g$: Generated Images $n\times C\times H\times W$ $\sigma^{2}(f)$ Variance per pixel (real) $\sigma^{2}(g)$ Variance per pixel (gen) $|\sigma^{2}(f)-\sigma^{2}(g)|$ Pixel-wise absolute difference $\mathcal{L}_{var}$ Average over all pixels
Figure 4: Computation of the proposed variance loss. After computing batch-wise pixel variance for real and generated tensors, the absolute difference is calculated at each pixel location, followed by averaging to obtain the final variance loss.
5 Experimental Setup
To evaluate the effectiveness of our proposed variance-based loss enhancement in pyramid pix2pix, we conduct experiments on a publicly available dataset (liu2022bci) of paired hematoxylin and eosin (H&E) and Immunohistochemistry (IHC) stained images. The dataset comprises high-resolution histopathological images from multiple tissue samples, ensuring diversity in staining patterns and structural variations. We preprocess all images by normalizing intensity values and resizing them to a fixed resolution suitable for deep learning models. The training and validation splits are maintained at an 80:20 ratio to ensure a balanced evaluation. Our experiments are implemented using PyTorch and executed on Google Colab, leveraging its Tesla T4 GPU acceleration. The model is trained with a batch size of 8, an initial learning rate of 0.0002, and the Adam optimizer to ensure stable convergence.
5.1 Dataset Analysis
The BCI dataset curated by (liu2022bci) serves as a comprehensive resource specifically designed for histopathological image translation tasks. It comprises 4873 paired image patches extracted from whole slide images of 51 patients, covering a broad spectrum of histological variations. The dataset includes all four categories of HER2 expression in IHC images: 0, 1+, 2+, and 3+. Each patch has a high resolution of 1024 x 1024 pixels, preserving fine-grained histopathological details necessary for accurate image-to-image translation. The paired nature of the dataset makes it particularly suitable for supervised learning approaches, where each H&E-stained image is directly associated with a corresponding IHC counterpart, as illustrated in Figure 5.
Figure 5: Representative image patches from the BCI dataset (liu2022bci), showcasing hematoxylin and eosin (H&E) stained images alongside their corresponding immunohistochemistry (IHC) images. The dataset provides paired samples for deep learning-based image translation and biomarker assessment.
5.14%24.7%45.9%24.3%IHC 0 (5.14%)IHC 1+ (24.7%)IHC 2+ (45.9%)IHC 3+ (24.3%)
@@ 5.1 Dataset Analysis
The dataset exhibits a non-uniform distribution of IHC image patches across the four HER2 expression levels, reflecting real-world prevalence and variability in HER2 staining patterns. As depicted in Figure 6, a higher proportion of images correspond to IHC 2+ expression, which represents a critical threshold category that requires additional FISH confirmation in clinical settings. The diversity of samples within the dataset is helpful in training models that can generalize effectively across different staining intensities and cellular morphologies. The variation in sample representation highlights the necessity of developing robust translation models capable of accurately generating high-quality IHC images, even for underrepresented categories.
IHC 0IHC 1+IHC 2+IHC 3+160160180180200200198198176176173173160160191191167167165165154154202202183183180180170170Average Color Means (RGB)R MeanG MeanB Mean
@@
IHC 0IHC 1+IHC 2+IHC 3+200200210210220220230230227.93227.93216.2216.2208.24208.24205.61205.61Average Dynamic Range
@@
A detailed analysis of the dataset reveals significant variability in terms of color distribution, sharpness, and dynamic range across different HER2 expression levels. Figure 7 illustrates differences in the average color means, highlighting distinct staining properties among different categories. Similarly, Figure 8 and Figure 9 provide insights into the texture and contrast variations between H&E and IHC images. These differences present challenges for existing generative models, particularly in maintaining structural integrity while ensuring realistic color translation. Addressing these variations remains essential for improving the fidelity of generated IHC images and enhancing their applicability in clinical diagnostic workflows.
5.2 Evaluation Metrics
To assess the quality of the generated immunohistochemistry (IHC) images from hematoxylin and eosin (H&E) stained images, we employ three widely used evaluation metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM) (wang2004image), and Fréchet Inception Distance (FID) (heusel2017gans). Each of these metrics provides a complementary perspective on the accuracy and perceptual quality of the translated images.
5.2.1 Peak Signal-to-Noise Ratio (PSNR)
@@
5.2.2 Structural Similarity Index (SSIM)
@@
$\mu_{x}$ $\text{ is the mean intensity of image }x,$
$\mu_{y}$ $\text{ is the mean intensity of image }y,$
$\sigma_{x}^{2}$ $\text{ is the variance of image }x,$
$\sigma_{y}^{2}$ $\text{ is the variance of image }y,$
$\sigma_{xy}$ $\text{ is the covariance between images }x\text{ and }y,$
@@
5.2.3 Fréchet Inception Distance (FID)
FID assesses the quality and diversity of the generated images by comparing the statistical distribution of deep features extracted from a pre-trained Inception v3 model (szegedy2016rethinking). Unlike PSNR and SSIM, which evaluate pixel-wise differences, FID provides a perceptual similarity measure at a higher abstraction level. It is defined as:
[…]
@@ 5.3 Experimental Framework
The BCI dataset (liu2022bci) we used consists of 4873 paired hematoxylin and eosin (H&E) and HER2 Immunohistochemistry (IHC) stained images. The IHC images are further classified into four categories based on staining intensity and membrane completeness, which help determine HER2 expression levels in tumor cells. IHC 0 is considered negative, showing no staining or faint, incomplete membrane staining in less than 10% of tumor cells. IHC 1+ is also negative but presents weak and incomplete membrane staining in more than 10% of tumor cells. IHC 2+ is considered equivocal, displaying weak to moderate complete membrane staining in over 10% of tumor cells, often requiring fluorescence in situ hybridization (FISH) for confirmation. IHC 3+, classified as HER2-positive, exhibits strong, circumferential membrane staining in more than 10% of tumor cells, confirming HER2 overexpression without the need for additional testing.
@@
The training dataset comprises images exhibiting diverse characteristics based on HER2 concentration levels, as reflected by variations in statistical properties such as color distribution, sharpness, and dynamic range. These variations are quantitatively illustrated in Figures 7, 8, and 9.
@@
To evaluate our proposed method, we compare the performance metrics, Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Fréchet Inception Distance (FID), against two baseline methods: pix2pix (isola2017image) and pyramid pix2pix (liu2022bci). To optimize computational efficiency, input images are cropped before being fed into the generator.
@@
The generator in our proposed framework utilizes a ResNet-9blocks architecture, while the discriminator follows the default PatchGAN configuration. Training is conducted with a batch size of 8 for 100 epochs. The optimization is performed using the Adam optimizer, with an initial learning rate of 0.0002 for the first 50 epochs, which then linearly decays to zero over the remaining epochs. The momentum parameter for Adam is set to 0.5. The generator is trained using the loss function defined in Equation 11, incorporating our proposed variance-based loss term.
6 Results
@@
To comprehensively assess the effectiveness of our proposed model, we evaluated its performance across multiple classes of images, focusing on its ability to generate high-fidelity immunohistochemistry (IHC) images from hematoxylin and eosin (H&E) stained images. Our results are presented in comparison with state-of-the-art models, including pix2pix (isola2017image) and pyramid pix2pix (liu2022bci), using the Breast Cancer Immunohistochemistry (BCI) dataset.
6.1 Comparison with Existing Methods on BCI Dataset
To assess the effectiveness of our proposed model in translating hematoxylin and eosin (H&E) stained images into high-fidelity immunohistochemistry (IHC) images, particularly for HER2-positive (IHC 3+) cases, we performed a detailed quantitative and qualitative analysis.
6.1.1 Quantitative Evaluation on Test Set
@@
Figure 10 illustrates the peak signal-to-noise ratio (PSNR), the structural similarity index (SSIM), and the Fréchet inception distance (FID) for different levels of expression of HER2. Our model consistently demonstrates superior performance, as evident by higher SSIM, PSNR, and lower FID scores, particularly for IHC 3+ images. This confirms that our method enhances the fidelity of IHC image synthesis, addressing the primary challenge of accurately reconstructing HER2-positive cases, which has remained a limitation of previous models.
IHC 0IHC 1+IHC 2+IHC 3+18182020222219.6719.6722.2222.2220.920.919.219.219.219.221.0721.0720.4520.4519.419.420.4820.4821.6521.6521.3421.3419.519.5Image ClassPSNRpix2pixpyramid pix2pixours
(a) PSNR
IHC 0IHC 1+IHC 2+IHC 3+0.30.30.40.40.50.50.520.520.470.470.440.440.390.390.460.460.440.440.420.420.380.380.560.560.480.480.460.460.420.42Image ClassSSIMpix2pixpyramid pix2pixours
(b) SSIM
IHC 0IHC 1+IHC 2+IHC 3+400400600600800800914914605605474474733733921921643643527527759759890890472472395395508508Image ClassFIDpix2pixpyramid pix2pixours
(c) FID
Figure 10: Class-wise quantitative comparison on the test set for Pix2pix, PyramidPix2pix, and the proposed method across IHC score categories (0, 1+, 2+, 3+). (a) PSNR and (b) SSIM (higher is better) measure pixel-level fidelity and structural similarity, while (c) FID (lower is better) reflects distribution-level realism. The proposed method achieves consistently improved PSNR/SSIM and lower FID across most classes, indicating better reconstruction quality and perceptual realism.
The evaluation metrics for our methodology, using the BCI dataset, are presented in comparison with existing approaches, namely pix2pix (isola2017image) and pyramid pix2pix (liu2022bci) in Table 5 and Figure 11. Notably, our method achieves the highest PSNR and SSIM scores, indicating superior pixel-wise fidelity and enhanced preservation of structural information, both essential for clinical interpretability. Moreover, the substantial reduction in FID highlights the strong resemblance between real and generated images in the feature space, underscoring the effectiveness of our model in achieving high-quality image translation.
Furthermore, our model outperforms existing models by reporting the best evaluation metrics for all classes of HER2 expression as presented in Tables 1-4 and Figure 11. This consistently superior performance across all classes highlights the robustness and reliability of our method.
@@
All these results indicate that our variance-based loss function plays a crucial role in mitigating mode collapse, resulting in a broader diversity of generated images and ensuring that the synthetic IHC images accurately capture the inherent variability observed in real-world HER2-positive cases. This advancement is critical for reliable computational pathology, where the preservation of fine-grained details in IHC images is paramount for clinical decision-making.
pix2pixpyramid pix2pixours1818202022222424262620.7420.7421.1521.1522.1622.16PSNR
(a) PSNR
pix2pixpyramid pix2pixours0.30.30.40.40.50.50.440.440.430.430.470.47SSIM
(b) SSIM
pix2pixpyramid pix2pixours400400600600800800472.6472.6516.75516.75346.37346.37FID
(c) FID
Figure 11: Comparison of image quality metrics across models. (a) PSNR, (b) SSIM, and (c) FID comparison between pix2pix, pyramid pix2pix, and the proposed method for the entire dataset representing all classes of HER2 expression.
6.1.2 Qualitative Comparison
@@
Figure 12: Comparison of different types of IHC images generated using different methods. The first column shows three randomly chosen input H&E images, the second column shows the corresponding real IHC images. The third, fourth and fifth columns represent IHC images generated using pix2pix, pyramid pix2pix and our method respectively. The images generated by our method are closer to the real IHC images.
6.1.3 Training Dynamics and Convergence Analysis
We compared the proposed variance-regularized method with Pix2pix and PyramidPix2pix on the validation set using PSNR and SSIM over 100 training epochs. As shown in Figure 13, the proposed method consistently achieves higher PSNR values than both baselines throughout training. The improvement remains stable across epochs, with a clear margin over Pix2pix and PyramidPix2pix. In addition, the PSNR curve of our method exhibits smoother and more stable convergence behavior. The higher PSNR indicates improved pixel-level reconstruction accuracy. By enforcing variance consistency between real and generated images at each spatial location, the model better preserves intensity dispersion and avoids excessive smoothing. Figure X(b) presents the SSIM comparison. The proposed method maintains the highest SSIM values for most of the training process and achieves superior structural similarity at convergence.
(a) Validation PSNR
(b) Validation SSIM
Figure 13: Validation performance comparison of the proposed method, PyramidPix2pix, and Pix2pix over 100 training epochs. Exponential moving average (EMA, $\alpha=0.05$) is applied to highlight convergence trends. The proposed method consistently achieves superior PSNR and SSIM compared to the baseline models.
A slight decrease in SSIM is observed during early epochs. This behavior is expected, as the variance regularization increases local intensity fluctuations while the generator adapts to match real-image variance statistics. After this adaptation phase, SSIM stabilizes and remains consistently higher than the baseline methods.
Figure 14 compares different training losses for our method against pyramidPix2pix and pix2pix. The increased generator loss magnitude arises from the additional variance regularization term and does not indicate inferior adversarial performance. Instead, its smooth convergence demonstrates stable multi-objective optimization.The discriminator loss remains stable and bounded throughout training, indicating balanced adversarial optimization without discriminator saturation or collapse.The gradual increase in adversarial loss reflects a strengthening discriminator and a more challenging adversarial game, rather than degradation in generated image quality. This is supported by simultaneous improvements in PSNR and SSIM.Although Pix2pix achieves lower L1 loss, this primarily reflects stronger pixel-wise averaging. The proposed method achieves superior PSNR and SSIM despite a higher L1 term, indicating improved structural and statistical fidelity rather than over-smoothed reconstruction.
(a) Generator Loss
(b) Adversarial Loss
(c) Discriminator Loss
(d) L1 Loss
Figure 14: Training loss comparison across the proposed method, Pix2pix, and PyramidPix2pix over 100 epochs. (a) Total generator loss, (b) adversarial loss, (c) discriminator loss, and (d) L1 reconstruction loss. The proposed method incorporates an additional variance-based regularization term, resulting in higher generator objective magnitude while maintaining stable discriminator dynamics and balanced adversarial training.
Figure 15 shows the mean variance loss per epoch we proposed in our method. The steady downward trend and absence of oscillatory instability demonstrate that the variance-based regularization integrates smoothly into the adversarial framework and converges reliably without destabilizing training.
Figure 15: Convergence of the variance regularization loss over 100 training epochs, showing stable and consistent optimization.
6.2 Generalizability: Evaluation on LLVIP Dataset
To evaluate the broader applicability of our model beyond medical imaging, we tested it on the facades dataset (isola2017image). The facades dataset is a commonly used benchmark in image-to-image translation tasks. It contains paired images of the façades of architectural buildings and their corresponding label maps. The dataset consists of approximately 506 image pairs (400 for training and 106 for testing). Each image pair typically consists of a label map or semantic segmentation showing structural elements and a target image which is a real photograph of the corresponding building façade.
Table 6 reports the evaluation metrics, demonstrating that our model achieves superior performance compared to Pyramid pix2pix in this domain. The ability of our model to generalize effectively to non-medical image-to-image translation tasks highlights its robustness and adaptability, further validating the impact of our variance-based loss function in mitigating mode collapse across different domains.
7 Discussion